# Coupling Brownian loop soups and random walk loop soups at all polynomial scales

### v1

@@ 3 Proof of the main theorem
We first treat the general case of continuous-time random walk loop soups on $\mathbb{Z}^{d},d\geq 1$, and then explain in Section 3.4 how to adapt this proof to the special case of discrete random walk loop soups on $\mathbb{Z}$ or $\mathbb{Z}^{2}$.
@@ 3.1 Preparation
In this subsection, we define some auxiliary random processes and loops, and construct a family of the random walk loop soups $\{\widetilde{\mathcal{A}}_{\lambda}\}_{\lambda>0}$, all on the same probability space $(\Omega,\mathcal{F},\mathbb{P})$. We will only look at continuous-time random walk loops, and will omit the word “continuous-time”.
@@
$\mathbb{P}\bigg[\sup_{0\leq s\leq 1}\big|\widetilde{\ell}(n,z;m)-\big(\widetilde{T}(n,z;m)/d\big)^{1/2}\ell(n,z,m)\big|\geq c(\eta,d)\log n\bigg]\leq c_{0}(\eta,d)n^{-\eta}.$ (3.1)
@@ 3.2 Construction of the Brownian loop soup
Recall the Poisson processes $N(n,z;t)$ and the coupled loops $\widetilde{\ell}(n,z;m),\ell(n,z;m)$ defined in Section 3.1. We further define independent complex random variables $Y(n,z;m)$ uniformely distributed on the hypercube $\big\{\sum_{j=1}^{d}x_{j}e_{j},|x_{j}|\leq 1/2\big\}$, and independent real random variables $T(n,z;m)$ with density
@@
3.3 Proof of Theorem 1.2 for continuous-time random walk loop soups
@@
$\mathbb{P}\left[N(n,z;\lambda)>0\text{ for some }n\geq N^{k},|z|\leq rN\right]\leq c_{1}\lambda r^{d}N^{(1-k/2)d}=c_{1}\lambda r^{2}N^{-a},$ (3.7)
@@
where the second inequality follows from the fact that $\gamma_{n,z,m}(sT)-Y(n,z;m)$ is a $d$-dimensional Bronwnian bridge with time length $T\sim 2n$, and that the maximum $M$ of a one-dimensional Brownian bridge with time length $1$ satisfies $\mathbb{P}(M>x)=\exp(-2x^{2})$.
@@
3.4 Proof of Theorem 1.2 for discrete-time random walk loop soups
The proof for discrete-time random walk loop soups on $\mathbb{Z}^{d}$ with $d=1,2$ is similar, so we will only outline the differences.
First, we do not need to add the loop soups $\{\widetilde{\mathcal{L}}_{\lambda}\}_{\lambda>0}$ as in Section 3.1, since there are by definition no loop with time length less than $2$. For integers $n\geq 1$ and $z\in\mathbb{Z}^{d}$, we similarly define independent Poisson processes $N(n,z;t)$, but with parameter $\widetilde{q}_{d}(n)$ given by (2.2) and (1.5). For each $n\geq 1$, $m\geq 1,z\in\mathbb{Z}^{d}$, let $\widetilde{\ell}(n,z;m)$ be an independent random walk bridge (loop) on $\mathbb{Z}^{d}$ with time length $2n$. Then the collection of loops
@@
Due to Theorem 2.2 and Lemma 2.3, we can also construct on the same probability space $(\Omega,\mathcal{F},\mathbb{P})$, i.i.d. Brownian loops $\ell(n,z;m),n\geq 1,z\in\mathbb{Z}^{d},m\geq 1$, distributed as $d$-dimensional Brownian bridges started at 0 with time length $1$, so that the following holds. By Lemma 2.5, there exist constants $c_{0}(\eta,d)$ and $c(\eta,d)$, such that for all $n\geq 1,z\in\mathbb{Z}^{d},m\geq 1$,
$\mathbb{P}\bigg[\sup_{0\leq s\leq 1}\big|\widetilde{\ell}(n,z;m)-(2n/d)^{1/2}\ell(n,z,m)\big|\geq c(\eta,d)\log n\bigg]\leq c_{0}(\eta,d)n^{-\eta}.$ (3.12)
@@ Acknowledgments
We thank Yifan Gao and Pierre Nolin for useful discussions. We also acknowledge the support by a GRF grant from the Research Grants Council of the Hong Kong SAR (project 11305825).

### v2

@@ 3 Proof of the main theorems
We first treat, in Sections 3.1, 3.2 and 3.3, the case of continuous-time random walk loop soups. We omit the adjective “continuous-time” in our description of random walk loop soups in these sections. We then explain, in Section 3.4, how to adapt this proof to the case of discrete-time random walk loop soups.
@@ 3.1 Preparation
In this subsection, we define some auxiliary random processes and loops, and construct a family of the random walk loop soups $\{\widetilde{\mathcal{A}}_{\lambda}\}_{\lambda>0}$, all on the same probability space $(\Omega,\mathcal{F},\mathbb{P})$.
@@
$\mathbb{P}\bigg[\sup_{0\leq s\leq 1}\big|\widetilde{\ell}(n,z;m)(sT)-\big(\widetilde{T}/d\big)^{1/2}\ell(n,z;m)(s)\big|\geq c(\eta,d)\log n\bigg]\leq c_{0}(\eta,d)n^{-\eta},$ (3.1)
where we write $T=T(n,z;m)$ and $\widetilde{T}=\widetilde{T}(n,z;m)$ for brevity.
@@ 3.2 Construction of the Brownian loop soup
Recall the Poisson processes $N(n,z;t)$ and the coupled loops $\widetilde{\ell}(n,z;m),\ell(n,z;m)$ defined in Section 3.1. We further define independent complex random variables $Y(n,z;m)$ uniformly distributed on the hypercube $\big\{\sum_{j=1}^{d}x_{j}e_{j},|x_{j}|\leq 1/2\big\}$, and independent real random variables $T(n,z;m)$ with density
@@
3.3 Proof of Theorem 1.2
@@
$\mathbb{P}\left[N(n,z;\lambda)>0\text{ for some }n\geq N^{k},|z|\leq rN\right]\leq c_{1}\lambda r^{d}N^{(1-k/2)d}=c_{1}\lambda r^{d}N^{-a},$ (3.7)
@@
where the second inequality follows from the fact that $\gamma_{n,z,m}(sT)-Y(n,z;m)$ is a $d$-dimensional Brownian bridge with time length $T\sim 2n$, and that the maximum $M$ of a one-dimensional Brownian bridge with time length $1$ satisfies $\mathbb{P}(M>x)=\exp(-2x^{2})$.
@@
3.4 Proof of Theorem 1.3
The proof for discrete-time random walk loop soups is similar, so we will only outline the differences. For simplicity, we keep the same notations as for the continuous-time random walk loop soups, but update some of their definitions.
First, we do not need to add the loop soups $\{\widetilde{\mathcal{L}}_{\lambda}\}_{\lambda>0}$ as in Section 3.1, since there are by definition no loop with time length less than $2$ in the discrete-time random walk loop soup. For integers $n\geq 1$ and $z\in\mathbb{Z}^{d}$, we similarly define independent Poisson processes $N(n,z;t)$, but with parameter $\widetilde{q}_{d}(n)$ given by (2.2), instead of $Q_{d}(n)$. For each $n\geq 1$, $m\geq 1,z\in\mathbb{Z}^{d}$, let $\widetilde{\ell}(n,z;m)$ be an independent random walk bridge (loop) on $\mathbb{Z}^{d}$ with time length $2n$. Then the collection of loops
@@
For $d=1,2$, due to Theorem 2.1 and Lemma 2.2, we can also construct on the same probability space $(\Omega,\mathcal{F},\mathbb{P})$, i.i.d. Brownian loops $\ell(n,z;m),n\geq 1,z\in\mathbb{Z}^{d},m\geq 1$, distributed as $d$-dimensional Brownian bridges started at 0 with time length $1$, so that the following holds.
$\mathbb{P}\bigg[\sup_{0\leq s\leq 1}\big|\widetilde{\ell}(n,z;m)(2ns)-(2n/d)^{1/2}\ell(n,z,m)(s)\big|\geq c(\eta,d)\log n\bigg]\leq c(\eta,d)n^{-\eta}.$ (3.12)
For $d\geq 3$, by Lemma 2.3, we get a slightly weaker bound, namely there exist $c(\eta,d)$, such that for all $n\geq 1,z\in\mathbb{Z}^{d},m\geq 1$,
$\mathbb{P}\bigg[\sup_{0\leq s\leq 1}\big|\widetilde{\ell}(n,z;m)(2ns)-(2n/d)^{1/2}\ell(n,z,m)(s)\big|\geq c(\eta,d)n^{1/4}\log n\bigg]\leq c(\eta,d)n^{-\eta}.$ (3.13)
The bounds (3.12) and (3.13) are the analogues of (3.1).
@@
For $d=1,2$, the rest of the proof is exactly the same as in Section 3.3. For $d\geq 3$, we still need to update the definitions of the events $A$, $A^{1}_{n,z,m}$ and $A^{2}_{n,z,m}$ to
$\begin{split}A:=\bigg\{\sup_{0\leq s\leq 1}&|\gamma_{n,z,m}(sT)-\widetilde{\gamma}_{n,z,m}(s\widetilde{T})|\geq 3k\,c(\eta,d)N^{k/4}\log N\\ &\text{ for some }|z|<rN,N^{\theta}<n<N^{k},1\leq m\leq N(n,z;\lambda)\bigg\}\end{split}$ (3.15)
and
$A^{1}_{n,z,m}=\bigg\{\sup_{0\leq s\leq 1}\big|(\widetilde{\gamma}_{n,z,m}(s\widetilde{T})-z)-(\widetilde{T}/d)^{1/2}T^{-1/2}(\gamma_{n,z,m}(sT)-Y(n,z;m))\big|\geq c(\eta,d)N^{k/4}\log N^{k}\bigg\},$
$A^{2}_{n,z,m}=\bigg\{\big((\widetilde{T}/d)^{1/2}T^{-1/2}-1\big)\sup_{0\leq s\leq 1}|\gamma_{n,z,m}(sT)-Y(n,z;m))|\geq c(\eta,d)N^{k/4}\log N^{k}\bigg\}.$
By (3.13), we deduce for all $N^{\theta}<n<N^{k}$, there exists a constant $c_{0}$ depending only on $a,\theta,d$, such that $\mathbb{P}(A^{1}_{n,z,m})\leq c_{0}N^{-a-d}$. We get a similar bound for $\mathbb{P}(A^{2}_{n,z,m})$, as in (3.10). The definition (3.15) of $A$ implies the bound (1.18) for the loops which are rescaled by $N^{-1}$ in space, noting that $k=2+2a/d$ by (3.6).
@@ Acknowledgments
We thank Yifan Gao, Xinyi Li and Pierre Nolin for useful discussions. We also acknowledge the support by a GRF grant from the Research Grants Council of the Hong Kong SAR (project 11305825).